# Towards Robust Spatio-Temporal Auto-Regressive Prediction: Adams-Bashforth Time Integration with Adaptive Multi-Step Rollout

### v1

@@ A.3 Contrast with Adams-Bashforth scheme
While RK methods offer superior accuracy in traditional numerical integration contexts, their fundamental requirements make them unsuitable for AR prediction tasks. The lack of ground truth for intermediate stages and the compound nature of prediction errors present critical challenges for practical implementation. These limitations explain the reason for our choice of the Adams-Bashforth scheme, which provides a more practical and robust approach to long-term AR prediction.
@@ Appendix B Evolution of the weights in AW3-based multi-step rollout
In Section 6, the weights for each loss term in the AW2-based multi-step rollout are visualized in Fig. 8. This section extends that analysis by presenting results from the AW3-based approach, using the same model configuration as in Fig. 8 but applying AW3 instead of AW2. Fig. 13 illustrates these results, revealing that after 1,000 epochs, the model begins to stabilize the weights of the first and last future steps, with the last time-step weight slightly exceeding that of the first. This indicates that AW3, which focuses only on the first and last time-step losses, still performs comparably well without the intermediate loss terms required by AW2—indeed, AW3 shows the MSE of 0.002 where AW2 shows 0.010 as in Table 3.

### v2

@@ A.3 Contrast with Adams-Bashforth scheme
A.4 Conclusion and implications for AR prediction
While RK methods offer high accuracy in traditional numerical integration, they inherently rely on intermediate states evaluated at fractional time steps that are typically unavailable in AR prediction datasets. Although end-to-end training could approximate RK behavior, this approach fundamentally undermines RK’s strength by combining four model outputs without truly reflecting the physically motivated multi-stage integration mechanism essential to these methods. This implementation adds unnecessary complexity without delivering RK’s accuracy benefits. Our work therefore prioritizes the Adams-Bashforth scheme as a more practical and robust alternative for long-term AR prediction, effectively leveraging available data while maintaining stability and computational efficiency for reliable extended time horizon predictions. Future research might explore hybrid approaches that successfully integrate multi-stage methods while preserving their physical principles within neural network architectures.
@@ Appendix B Evolution of the weights in AW3-based multi-step rollout
In Section 6, the weights for each loss term in the AW2-based multi-step rollout are visualized in Fig. 9. This section extends that analysis by presenting results from the AW3-based approach, using the same model configuration as in Fig. 9 but applying AW3 instead of AW2. Fig. 14 illustrates these results, revealing that after 1,000 epochs, the model begins to stabilize the weights of the first and last future steps, with the last time-step weight slightly exceeding that of the first. This suggests that AW3, which initially considers both the first and last time-step losses, remains highly effective by gradually shifting its focus to the last time step, ensuring greater accuracy in long-term rollouts.